### v2

# Understanding Higher-Order Correlations Among Semantic Components in Embeddings

Abstract
Independent Component Analysis (ICA) offers interpretable semantic components of embeddings. While ICA theory assumes that embeddings can be linearly decomposed into independent components, real-world data often do not satisfy this assumption. Consequently, non-independencies remain between the estimated components, which ICA cannot eliminate. We quantified these non-independencies using higher-order correlations and demonstrated that when the higher-order correlation between two components is large, it indicates a strong semantic association between them, along with many words sharing common meanings with both components. The entire structure of non-independencies was visualized using a maximum spanning tree of semantic components. These findings provide deeper insights into embeddings through ICA.
Understanding Higher-Order Correlations Among
Semantic Components in Embeddings
Momose Oyama1,2 Hiroaki Yamagiwa1 Hidetoshi Shimodaira1,2 1Kyoto University 2RIKEN oyama.momose@sys.i.kyoto-u.ac.jp, hiroaki.yamagiwa@sys.i.kyoto-u.ac.jp, shimo@i.kyoto-u.ac.jp
1 Introduction
Figure 1: Heatmap visualization of 300-dimensional SGNS embeddings transformed by PCA and ICA, with axes sorted by variance and skewness, respectively. Each embedding has been normalized to have a norm of 1 for better visual interpretation. For each axis, the top 4 words (frequency $n_{w}\geq 100$ in text8) with largest component values were used. The first 100 axes are displayed in the top panels, and the first 5 axes with the word labels are displayed in the bottom panels. See Appendices A, B and G for details.
Embeddings play an important role in natural language processing, ranging from word embeddings Mikolov et al. (2013) to internal representations in language models Devlin et al. (2019); Brown et al. (2020); Touvron et al. (2023). Understanding how embeddings represent meaning is crucial for unraveling black box NLP models.
Independent Component Analysis (ICA) Hyvärinen and Oja (2000) is an effective method for visualizing and interpreting the geometric structure of embeddings Musil and Mareček (2024); Yamagiwa et al. (2023). Just as PCA aims to make coordinate axes uncorrelated, ICA seeks to transform the coordinate axes into statistically independent components. The resulting axes from ICA tend to have sparser component values with a few larger values compared to PCA, which increases interpretability as the axes can be seen as specific semantic components (Fig. 1).
However it has been pointed out that the estimated ‘independent components’ are only approximately independent Hyvärinen et al. (2001); Sasaki et al. (2013, 2014). This is because many real-world datasets cannot be accurately represented as a linear combination of independent components, contradicting the assumption of ICA theory.
In this study, we aim to further interpret the results of applying ICA to embeddings by focusing on the non-independence between ‘independent components’. We measure the degree of non-independence by calculating higher-order correlations between components and find that components with large higher-order correlations can be interpreted as having strong semantic associations. The entire structure is revealed by visualizing the maximum spanning tree of semantic components with higher-order correlations as edge weights.
2 Review: ICA-Transformed Embeddings
Procedure of ICA.
For a centered embedding matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ that represents the meanings of $n$ words by $d$-dimensional vectors, ICA [^1] For the computation, FastICA Hyvärinen (1999) implemented in scikit-learn Pedregosa et al. (2011) is used. seeks a transformation $\mathbf{S}=\mathbf{X}\mathbf{B}$ such that each component $S_{1},\cdots,S_{d}$ of the transformed matrix $\mathbf{S}=[S_{1},\cdots,S_{d}]$ is as statistically independent as possible [^2] The $k$-th component $S_{k}$ is also referred to as Axis $k$.. The transformation $\mathbf{B}$ can be expressed as the product of the whitening transformation matrix $\mathbf{A}$ (e.g., PCA transformation) and the orthogonal transformation matrix $\mathbf{R}_{\mathrm{ica}}$, i.e., the resulting $\mathbf{S}$ is represented as
$\mathbf{S}=\mathbf{X}\mathbf{A}\mathbf{R}_{\mathrm{ica}}.$ (1)
[…]
Comparison of PCA and ICA.
Figure 2: Scatterplots of normalized word embeddings along the 10th and 20th axes. The axes for PCA and ICA-transformed embeddings were arranged in descending order of variance and skewness, respectively. In both transformations, the components are uncorrelated.
Figure 2 shows that ICA can find the ‘spiky and interpretable shape’ of the embedding distribution (e.g., “biology” and “stars” for the 10th and 20th axes, respectively), but PCA cannot. This is because ICA determines the coordinate axes toward high non-gaussianity, while PCA only considers variance information.
3 Higher-Order Correlations Among Estimated Independent Components
Non-Independence in Real-World Data.
The ‘independent components’ estimated by ICA on real-world data are uncorrelated but not completely independent, with dependencies existing between components Hyvärinen et al. (2001); Sasaki et al. (2013, 2014). This is because ICA assumes a linear decomposition into independent components, an assumption frequently violated in reality.
Higher-Order Correlation.
To quantify non-independencies, methods like mutual information and Hilbert-Schmidt Independence Criterion (HSIC) Gretton et al. (2005) exist. Here we use the higher-order correlation, the simplest measure in terms of computation and formulation. This measure is expressed as follows:
$\mathrm{E}(S_{i}^{2}S_{j}^{2})=\frac{1}{n}\sum_{t=1}^{n}\mathbf{S}_{t,i}^{2}\mathbf{S}_{t,j}^{2}.$ (2)
Here, $\mathbf{S}$ is the whitened matrix [^5] The components are (i) centered: $\mathrm{E}(S_{i})=0$, the mean of each component is 00, (ii) scaled: $\mathrm{E}(S_{i}^{2})=1$, the variance of each component is $1$, and (iii) uncorrelated: $\mathrm{E}(S_{i}S_{j})=0$, the correlations are all zero.. This can also be interpreted as the covariance between $S_{i}^{2}$ and $S_{j}^{2}$, plus one, as $\mathrm{cov}(S_{i}^{2},S_{j}^{2})=\mathrm{E}((S_{i}^{2}-1)(S_{j}^{2}-1))=\mathrm{E}(S_{i}^{2}S_{j}^{2})-1$. If $S_{i}$ and $S_{j}$ are independent of each other, then $\mathrm{E}(S_{i}^{2}S_{j}^{2})=\mathrm{E}(S_{i}^{2})\mathrm{E}(S_{j}^{2})=1$. Thus, the deviation of $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ from $1$ is the degree of dependence between $S_{i}$ and $S_{j}$.
Figure 3: Heatmaps of the correlation coefficient $\mathrm{E}(S_{i}S_{j})$ and the higher-order correlation $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ of component pairs $(S_{i},S_{j})$ from ICA on 300-dimensional SGNS embeddings. See Appendix C for details.
Figure 3 shows that the estimated independent components of the embeddings are uncorrelated but not completely independent, with varying degrees of higher-order correlations across pairs. These $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ values provide a useful metric of association, as demonstrated in the following section.
$\mathrm{E}(S_{0}^{2}S_{82}^{2})=1.927$ $\mathrm{E}(S_{6}^{2}S_{96}^{2})=2.032$ $\mathrm{E}(S_{12}^{2}S_{66}^{2})=1.975$ $\mathrm{E}(S_{16}^{2}S_{118}^{2})=2.124$ $\mathrm{E}(S_{56}^{2}S_{126}^{2})=1.861$ $\mathrm{E}(S_{63}^{2}S_{210}^{2})=2.964$ Axis 0 Axis 82 Axis 6 Axis 96 Axis 12 Axis 66 Axis 16 Axis 118 Axis 56 Axis 126 Axis 63 Axis 210 dishes beer el o rabbi judah blood disorder cpu pointer organization unesco sauce beers spanish portuguese talmud israelites organs mental microprocessor return international itu fried ale nacional paulo rabbis yahweh liver disorders processor string organizations interpol dish brewing jos rio torah elisha kidney symptoms cpus pointers interpol observer cooked yeast de portugal jewish isaiah tissue bipolar intel node standardization temporary $\mathrm{E}(S_{0}^{2}S_{23}^{2})=0.990$ $\mathrm{E}(S_{6}^{2}S_{13}^{2})=0.992$ $\mathrm{E}(S_{12}^{2}S_{57}^{2})=0.993$ $\mathrm{E}(S_{16}^{2}S_{57}^{2})=0.996$ $\mathrm{E}(S_{56}^{2}S_{197}^{2})=0.982$ $\mathrm{E}(S_{63}^{2}S_{18}^{2})=1.073$ Axis 0 Axis 23 Axis 6 Axis 13 Axis 12 Axis 57 Axis 16 Axis 57 Axis 56 Axis 197 Axis 63 Axis 18 dishes statesman el windows rabbi s blood s cpu population organization actress sauce astronomer spanish os talmud and organs and microprocessor median international footballer fried philosopher nacional unix rabbis was liver was processor estimated organizations musician dish johann jos linux torah in kidney in cpus residing interpol actor cooked mathematician de microsoft jewish by tissue by intel total standardization singer
Table 1: (Top Row) Six randomly selected pairs of components from the top 50 pairs with the highest $|\mathrm{E}(S_{i}^{2}S_{j}^{2})-1|$ values. For each component, the top 5 words (frequency $n_{w}\geq 100$ in text8) with the largest component values are listed. (Bottom Row) Component pairs with small $\mathrm{E}(S_{i}^{2}S_{k}^{2})$ values. For each component $S_{i}$ with the smaller axis number in the pairs $(S_{i},S_{j})$ in the top row, a component $S_{k}$ with the smallest value of $|\mathrm{E}(S_{i}^{2}S_{k}^{2})-1|$ was selected.
$k=1$ $k=2$ $k=3$ $k=4$ $k=5$ List-2 (top-$k$) 69.0 65.0 64.0 64.5 56.5 List-3 (bottom 30%) 27.0 33.0 32.5 33.5 40.5 Can’t decide 4.0 2.0 3.5 2.0 3.0
Table 2: The percentage of each list judged by the GPT model to be more semantically related to List-1.
$\mathrm{E}(S_{10}^{2}S_{2}^{2})=2.323$ $\mathrm{E}(S_{10}^{2}S_{16}^{2})=1.947$ $\mathrm{E}(S_{10}^{2}S_{160}^{2})=1.811$ $\mathrm{E}(S_{27}^{2}S_{11}^{2})=1.643$ $\mathrm{E}(S_{27}^{2}S_{64}^{2})=1.997$ $\mathrm{E}(S_{27}^{2}S_{104}^{2})=1.605$ Axis $10$ Axis $2$ Axis $10$ Axis $16$ Axis $10$ Axis $160$ Axis $27$ Axis $11$ Axis $27$ Axis $64$ Axis $27$ Axis $104$ dna acid dna blood dna evolution greek gaius greek goddess greek archaeological proteins hydrogen proteins organs proteins evolutionary greece caesar greece gods greece neolithic rna acids rna liver rna darwin athens augustus athens deity athens bc mrna oh mrna kidney mrna selection athenian lucius athenian deities athenian pottery $w_{k}$ $\mathbf{S}_{k,10}^{2}\mathbf{S}_{k,{2}}^{2}$ $w_{k}$ $\mathbf{S}_{k,10}^{2}\mathbf{S}_{k,{16}}^{2}$ $w_{k}$ $\mathbf{S}_{k,10}^{2}\mathbf{S}_{k,{160}}^{2}$ $w_{k}$ $\mathbf{S}_{k,27}^{2}\mathbf{S}_{k,{11}}^{2}$ $w_{k}$ $\mathbf{S}_{k,27}^{2}\mathbf{S}_{k,{64}}^{2}$ $w_{k}$ $\mathbf{S}_{k,27}^{2}\mathbf{S}_{k,{104}}^{2}$ ribose 3755.7 adenylate 2079.8 utr 2381.5 laertius 898.3 demeter 2348.5 tiryns 1690.6 deoxyribose 2963.9 effectors 1842.5 reticulum 1942.0 preveza 788.0 hephaestus 2204.3 knossos 1348.1 phosphodiester 2850.2 antisense 1639.9 genomic 1668.6 xanthippus 764.2 hestia 2021.5 mycenaean 1205.6 biosynthesis 2510.1 cyclase 1638.9 homozygous 1599.1 rhadamanthus 735.5 hera 1744.6 lendering 1124.7 methyltransferase 2482.9 myosin 1201.8 cleaved 1181.0 thracians 711.8 cronos 1720.2 hissarlik 1103.1 pyrimidine 2399.6 axons 1144.2 tubulin 1152.4 alexandri 705.2 aphrodite 1675.9 melos 1006.3
Table 3: For 6 component pairs $(S_{i},S_{j})$ selected from adjacent component pairs in the MST defined in Sec. 5, the top 6 words and their corresponding $\mathbf{S}_{t,i}^{2}\mathbf{S}_{t,j}^{2}$ values that contribute the most to the $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ value are shown.
4 Interpretation of Higher-Order Correlations as Semantic Relevance
4.1 Degree of Semantic Relevance
We show that the values of higher-order correlations $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ can be interpreted as the degree of associations between semantic components.
Results: Top Row of Table 1.
The meanings of each component, represented by the listed words in component pairs with high $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ values, are strongly related. For example, focusing on Axis 00 and Axis $82$, a pair with particularly large values of $\mathrm{E}(S_{i}^{2}S_{j}^{2})$, we can interpret that Axis 00 has a meaning associated with “dishes” and Axis $82$ with “beer”, suggesting that there is a semantic relationship between them.
Results: Bottom Row of Table 1.
On the other hand, for component pairs with $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ values close to $1$, indicating that the components are considered independent, there is no clear relevance between the meanings of the components. For example, looking at the pair of Axis 00 and Axis $23$, which has a small $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ value, we can interpret that Axis 00 represents “dishes” and Axis $23$ represents “polymath”, and there is no direct semantic relationship between them.
Detailed results are shown in Appendix G.
4.2 Quantitative Evaluation via GPT-4o mini
We conducted experiments to quantitatively evaluate whether higher-order correlations between ICA components represent semantic relationships.
Settings.
Our experimental procedure was as follows. We first selected the top 100 ICA components, ranked by skewness. For each component $i$ ($i=0,\ldots,99$), we created three word lists: Word list-1 comprised the top 5 words from component $i$, Word list-2 contained the top 5 words from the $k$-th most correlated component with component $i$ ($k=1,\ldots,5$), and Word list-3 consisted of the top 5 words from a randomly selected low-correlation component (chosen from the bottom 30% of correlated components). Using these lists, we generated pairs (list-1, list-2) and (list-1, list-3), and queried GPT-4o mini to determine which pair was more semantically related [^6] To mitigate potential biases, we randomly shuffled the order of the lists in the pairs and repeated this process with the reversed order: (list-1, list-3) and (list-1, list-2).. This procedure was executed for all 100 components, resulting in 200 total comparisons for each value of $k$ from 1 to 5. The specific prompt used for GPT-4o mini model is provided in Appendix D.
Results and Discussion.
Table 2 shows the result of the experiment. We can see that component pairs with higher-order correlations tend to be more semantically related (69.0% for $k=1$ vs 27.0% for bottom 30%), and that semantic relatedness gradually declines as correlation decreases (69.0% at $k=1$ to 56.5% at $k=5$). These results quantitatively demonstrate that higher-order correlations between ICA components effectively reflect semantic relatedness between corresponding words.
4.3 Decomposition of Semantic Relevance
Figure 4: Scatter plots of normalized word embeddings for axis pairs (10, 2) and (27, 64) with large values of higher-order correlations. Blue-labeled words are the top 4 words for each axis’s component values, while red-labeled words are the top 6 words for the values of $\mathbf{S}_{t,i}^{2}\mathbf{S}_{t,j}^{2}$. See Appendix C for all the pairs in Table 3.
For a component pair $(S_{i},S_{j})$, words $w_{t}$ with large values of $\mathbf{S}_{t,i}^{2}\mathbf{S}_{t,j}^{2}$ are considered to make a significant contribution to the higher-order correlation $\mathrm{E}(S_{i}^{2}S_{j}^{2})=\frac{1}{n}\sum_{t=1}^{n}\mathbf{S}_{t,i}^{2}\mathbf{S}_{t,j}^{2}$. Here we investigate words that significantly contribute to the $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ values and gain a more concrete understanding of the relationships between components.
Results.
Table 3 presents component pairs selected from the maximum spanning tree $T_{150}$ (Sec. 5) and words significantly contributing to their $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ values. These words often relate to the meanings of both components, demonstrating additive compositionality. For example, in the Axis 10 and Axis 2 pair, words like ribose, deoxyribose, phosphodiester, biosynthesis, methyltransferase, and pyrimidine notably contribute to the $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ value, linking biomolecules and chemical components. Detailed results are shown in Appendix G.
Visualization.
Figure 4 shows word embedding scatter plots for axis pairs (10, 2) and (27, 64) with large higher-order correlations to illustrate the distribution of words with significant contributions to higher-order correlations. Unlike a typical independent component scatter plot (Fig. 2), these exhibit many words with large component values in both axes, reflecting the meanings of both axes and demonstrating the additive compositionality of embeddings. For the (10, 2) pair, words that notably contribute to the $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ (ribose, deoxyribose, phosphodiester, biosynthesis, methyltransferase, and pyrimidine) appear with significant values in both components. This abundance of words sharing both semantic components is characteristic of pairs with large higher-order correlations. Detailed results are shown in Appendix C.
5 Visualization of Non-Independence Structure
Figure 5: Subtrees of MST $T_{150}$ defined in Sec. 5. Each node represents an independent component $S_{k}$ (i.e., Axis $k$) estimated by ICA. The label of each node is “$k$ : $\mathrm{TopWord}(k)$”, where $\mathrm{TopWord}(k)$ is the word with the largest component value along axis $k$ among words with frequency $n_{w}\geq 100$ in the text8 corpus. The color of the edge between nodes $(i,j)$ represents the magnitude of the $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ value between the components, with darker edge colors indicating larger values.
In this section, we construct a maximum spanning tree (MST) based on higher-order correlations to visualize the non-independence between estimated independent components.
Settings.
The 300 ICA components, originally ordered by skewness with $i=0,\cdots,299$, were re-sorted in descending order of semantic component consistency scores to prioritize axes that are more easily interpretable as specific semantic components. The semantic component consistency scores were determined by a word intrusion task Chang et al. (2009). A higher consistency score indicates easier interpretability. Details of the scoring methods are provided in Appendix E.1. We introduce the notation $\sigma$ to map the order of consistency scores to the original axis numbers in the skewness order: $\sigma(j)$ represents the axis number in the skewness sort for the axis with the $j$-th highest consistency score. Then, we consider a weighted complete graph $G_{150}$, with 150 components having high consistency scores $S_{\sigma(i)}~{}(i\in{0,\cdots,149})$ as nodes. For the edge between the node pair $(S_{i},S_{j})$, we set $c_{ij}=\mathrm{E}(S_{i}^{2}S_{j}^{2})$ as the weight. To visualize and interpret $G_{150}$, we compute the maximum spanning tree (MST)777 We used minimum_spanning_tree implemented in NetworkX Hagberg et al. (2008) for the computation of the MST. See Appendix E.2 for details. $T_{150}$, a spanning tree that maximizes the sum of $c_{ij}$ in the graph $G_{150}$. MST was relatively more interpretable than other subgraphs of graph $G_{150}$, providing a good balance between visibility and element relationships.
Interpretation of the MST.
The MST $T_{150}$ represents a graph structure expressing the non-independence between estimated independent components. Since the edges in $T_{150}$ connect component pairs with large higher-order correlations, we can interpret that there is a strong relationship between the components connected by these edges. Furthermore, the subtrees of $T_{150}$ represent groups of semantically related components, and the components within these groups tend to have similar meanings.
Results and Discussion.
[…]
6 Conclusion
Both ICA and PCA transformations make the components uncorrelated. ICA goes further by making the components nearly independent, but some non-independence still remains. In this study, we used higher-order correlations to quantify the non-independence between the components in the ICA-transformed word embeddings. By interpreting these as the semantic associations between the components and visualizing the overall structure, we can gain a deeper understanding of the latent semantic structure within the embeddings.
Limitations
  • The embeddings used in the experiments are limited to SGNS word embeddings. For a more thorough analysis, it is necessary to conduct experiments using various types of embeddings.
  • For large embedding matrices with a high number of data points $n$, ICA may fail to converge within a practical timeframe. To overcome this, we suggest using subsampled data to estimate the ICA transformation matrix, which can then be applied to unseen embedding vectors.
  • When $n$ is large, calculating higher-order correlations (eq. 2) may become computationally intensive. This calculation is similar to the computation of the variance-covariance matrix and can be approximated by subsampling data points. Further speedup can be achieved by parallelization of the computation.
  • Since ICA leverages the non-Gaussianity of embedding distributions, it is not suitable for analysis if the original embeddings follow a multivariate Gaussian distribution.
Ethics Statement
This study complies with the ACL Ethics Policy.
Acknowledgements
This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST BOOST JPMJBS2407, JST SPRING JPMJSP2110.
Code Availability
Code is available at https://github.com/momoseoyama/hoc.
Appendix A Details of Experimental Settings
The word embeddings used in the experiments were trained using Skip-gram with Negative Sampling (SGNS). The parameters used to train SGNS are summarized in Table 4. The corpus used for training is the text8 corpus Mahoney (2011), and the number of vocabulary words is $n=253{,}854$.
Dimensionality 300
Epochs 100
Window size $h$ 10
Negative samples $\nu$ 5
Learning rate 0.025
Min count 1
Table 4: SGNS parameters.
Appendix B Remarks on Axis $57$ in Figure 1
Figure 6: Boxplot of correlation coefficients between word frequency $n_{w}$ and the component values for the 00th to $99$th axes of the ICA-transformed embeddings. Axis $57$ shows a particularly high correlation coefficient.
Figure 7: Scatter plot of word frequency $n_{w}$ versus the component values of the $57$th axis of the ICA-transformed embeddings. Words used in Fig. 1 are highlighted in dark blue. The regression line and coefficient of determination were calculated for words with a frequency of $n_{w}\geq 10$.
Figure 8: Scatter plot of word frequency $n_{w}$ versus the component values of the 00th axis of the ICA-transformed embeddings. The settings are the same as in Fig. 7.
An interesting vertical streak is observed in Axis $57$ of the heatmap for ICA-transformed embeddings in Fig. 1. This streak can be explained by several factors. As shown in Fig. 6, Axis $57$ exhibits a strong correlation between component values and word frequencies $n_{w}$, suggesting that Axis 57 is more associated with word frequency than with a specific semantic meaning. Additionally, the words used in Fig. 1 were selected from those appearing more than 100 times in the text8 corpus, resulting in a bias towards high-frequency words. This tendency is further illustrated in Fig. 7, which demonstrates that words used in the heatmap $(n_{w}\geq 100)$ tend to have larger component values along Axis $57$. In contrast, Fig. 8 shows that for axes with weak correlation to word frequency, the words used in the heatmap do not exhibit notably large component values. Consequently, large component values were observed along Axis 57 in the heatmap, a pattern that was unique to Axis 57 and not observed in other axes.
(a) For $(i,j)$ pairs where $i<j$.
(b) For all $(i,j)$ pairs, including $i=j$.
Figure 9: Histograms of higher-order correlations.
Appendix C Higher-Order Correlations
C.1 Distribution of Higher-Order Correlations
Figures 9(a) and 9(b) show histograms of higher-order correlations $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ for pairs where $i<j$ and for all pairs including $\mathrm{E}(S_{i}^{4})$ where $i=j$, respectively. While there are component pairs where $\mathrm{E}(S_{i}^{2}S_{j}^{2})<1$, in Fig. 3, the range of values was truncated between $1.0$ and $2.5$ for visualization purposes.
C.2 Scatterplots for Independent Axes
Figure 10: Scatter plots of normalized word embeddings for axis pairs in Table 3. Blue-labeled words are the top 4 words for each axis’s component values, while red-labeled words are the top 6 words for values of $\mathbf{S}_{t,i}^{2}\mathbf{S}_{t,j}^{2}$.
Figure 11: Scatter plots of normalized word embeddings for 24 axis pairs. Blue-labeled words are the top 4 words for each axis’s component values, while red-labeled words are the top 6 words for values of $\mathbf{S}_{t,i}^{2}\mathbf{S}_{t,j}^{2}$.
Complementary Results for Sec. 4.3.
Table 3 in Sec. 4.3 presented words with significant contributions to higher-order correlations for six axis pairs. While the main text illustrated the distribution of these highly contributing words through scatter plots for the two selected pairs, Figure 10 provides scatter plots for all the six pairs.
The Relationship Between the Magnitude of Higher-Order Correlations and the Appearance of Scatter Plots.
Figure 11 presents the scatter plots of word embeddings for 24 component pairs, each with different higher-order correlation values. We can see that as the magnitude of higher-order correlation increases, the number of words with large component values along both axes increases as well. The selection of these 24 pairs was conducted as follows: First, we considered 150 components $S_{\sigma(0)},\cdots,S_{\sigma(149)}$ with high semantic consistency (see Appendix E.1 for the calculation method). We then sorted all possible component pairs $(S_{i},S_{j})$ $(i,j\in{\sigma(0),\cdots,\sigma(149)})$ based on the value of $|\mathrm{E}(S_{i}^{2}S_{j}^{2})-1|$. We established 24 equally spaced grids between the minimum and maximum values, and selected pairs closest to each grid point without repetition.
Words with Significant Contributions to Higher-Order Correlations.
We have observed words with significant contributions, i.e., with large values of $\mathbf{S}_{t,i}^{2}\mathbf{S}_{t,j}^{2}$, to higher-order correlations $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ in Table 3 in Sec. 4.3, and will see further examples in Tables 8 and 9 in Appendix G. Such words are labeled in red in the scatter plots in Fig. 4 in Section 4.3, as well as in Figs. 10 and 11 in Appendix C.2. For axis pairs with large higher-order correlations, we observe a large number of words that make significant contributions to the higher-order correlations. The meanings of these words include both axes’ meanings, demonstrating the additive compositionality of embeddings.
Appendix D Prompt Used for Evaluation by GPT-4o mini
The specific prompt used for the GPT-4o mini model is provided below.
Question:
You are given 2 list pairs (A, B), (C, D).
If one pair is more semantically relevant than the other, answer the pair.
If you cannot determine, answer "XX".
List pair (A, B): ({wordlist_1}, {wordlist_2})
List pair (C, D): ({wordlist_1}, {wordlist_3})
Output:
"AB" if (A, B) is more semantically related
"CD" if (C, D) is more semantically related
"XX" if equally related, or you cant decide
Respond with only AB, CD, or XX.
When conducting the experiment, we took the following steps to eliminate any potential biases arising from the order of word sequences and specific label names.
  1. 1.
    To remove the influence of word order within lists, we randomly shuffled the words in wordlist_1, wordlist_2, and wordlist_3.
  2. 2.
    To prevent bias in output labels, we conducted the experiment twice, swapping wordlist_2 and wordlist_3 between runs.
  3. 3.
    To account for order bias in the prompt, we randomly alternated the order of the following two lines:
    List pair (A, B): ({wordlist_1}, {wordlist_2})
    List pair (C, D): ({wordlist_1}, {wordlist_3})
Appendix E Details of Visualization of Non-Independence Structure
E.1 Scoring ICA Axes: Word Intrusion Task
We assigned a semantic coherence score to each axis of the ICA-transformed embeddings using the word intrusion task method Chang et al. (2009).
Word Intrusion Task.
[…]
Selection of the Intruder Word.
In order to select the intruder word for the set of top $k$ words of each axis $a\in\{1,\ldots,d\}$, denoted as $\mathrm{top}_{k}(a)$, we randomly chose a word from a pool of words that satisfy both of the following criteria simultaneously: (i) the word ranks in the lower $50\%$ in terms of the component value on the axis $a$, and (ii) it ranks in the top $10\%$ in terms of the component value on some axis other than $a$. For each axis, $L=100$ intruder words are randomly selected, and $W_{\mathrm{int}}(a)$ denotes the set of these $L$ intruder words.
Scoring Method.
For the consistency score of the meaning of each axis $a$, $\mathrm{Score}(a)$, we adopted the metric proposed by Sun et al. (2016).
$\mathrm{Score}(a)=\frac{\mathrm{InterDist}(a)}{\mathrm{IntraDist}(a)}$
$\mathrm{IntraDist}(a)$ $=\sum_{\begin{subarray}{c}w_{i},w_{j}\in\mathrm{top}_{k}(a)\\ w_{i}\neq w_{j}\end{subarray}}\frac{\mathrm{dist}(w_{i},w_{j})}{k(k-1)}$
$\mathrm{InterDist}(a)$ $=\underset{w\in W_{\mathrm{int}}(a)}{\mathrm{mean}}\sum_{w_{i}\in\mathrm{top}_{k}(a)}\frac{\mathrm{dist}(w_{i},w)}{k}$
In this formula, we defined $\mathrm{dist}(w_{i},w_{j})=\|\mathbf{s}_{i}-\mathbf{s}_{j}\|$ for the ICA-transformed embeddings. Here, $\mathrm{IntraDist}(a)$ denotes the average distance between the top $k$ words, and $\mathrm{InterDist}(a)$ represents the average distance between the top words and the intruder words. The score is higher when the intruder words are further away from the set $\mathrm{top}_{k}(a)$. Therefore, this score serves as a quantitative measure of the ability to identify the intruder word, thus it is used as a measure of the consistency of the meaning of the top $k$ words and the interpretability of axes.
E.2 Entire Visualization of MST
Figure 12 is the visualization of maximum spanning tree (MST) $T_{150}$ defined in Sec. 5. For a graph $G_{150}$, where the cost between nodes $i$ and $j$ defined as $c_{ij}=E(S_{i}^{2}S_{j}^{2})$, the algorithm to find the MST $T$ is a greedy method that maximizes the total sum of costs, $\sum_{(i,j)\in T}c_{ij}$, subject to $T$ being a spanning tree. The greedy algorithm selects edges in decreasing order of $c_{ij}$ while adhering to the tree constraint. Due to the monotonicity of $f(x)=1/x$, the decreasing order of $c_{ij}$ is equivalent to the increasing order of $1/c_{ij}$. Thus, computing the MST $T_{150}$ in the graph $G_{150}$ is equivalent to finding the minimum spanning tree, which minimizes the sum of $1/c_{ij}$.
Figure 12: Visualization of the entire MST $T_{150}$ defined in Sec. 5. Each color of the nodes represents one of the ten clusters obtained by Spectral Clustering.
Appendix F Dimensionality Reduction via MST
$d=2$ $d=5$ $d=10$ $d=20$ $d=50$ $d=100$ Random Clustering on components (PCA) 0.04 0.08 0.12 0.16 0.23 0.29 Random Clustering on components (ICA) 0.04 0.08 0.12 0.17 0.24 0.29 Spectral Clustering on MST (PCA) 0.03 0.08 0.13 0.17 0.24 0.30 Spectral Clustering on MST (ICA) 0.06 0.13 0.18 0.23 0.28 0.31
Table 5: Word similarity scores for dimensionality reduction.
We experimentally confirmed that the structure of higher-order correlations between components can be applied to the dimensionality reduction of embeddings. Specifically, we performed Spectral Clustering on the maximum Spanning Tree (MST) $T_{300}$, which was computed based on the higher-order correlations between components. By reducing the dimensionality of the embeddings through averaging the clustered axes, we verified that the accuracy degradation in Word Similarity Tasks was mitigated compared to random clustering.
Experimental Settings.
We conducted our experiments using the 300-dimensional word embeddings (SGNS). These embeddings were subjected to PCA and ICA to obtain components for clustering. We employed two clustering methods: (1) Random Clustering and (2) Spectral Clustering on the maximum Spanning Tree (MST) $T_{300}$, which was computed based on the higher-order correlations calculated using Eq. 2. Clustering was performed with the number of clusters ranging from 2 to 100. Dimensionality reduction was achieved by averaging the clustered axes, resulting in reduced dimensions from $d=2$ to $d=100$. The performance of the lower-dimensional embeddings was evaluated through Word Similarity Tasks. For the Word Similarity Tasks, we utilized six datasets: MEN Bruni et al. (2014), WS353 Finkelstein et al. (2002), MTurk Radinsky et al. (2011), RW Luong et al. (2013), SimLex999 Hill et al. (2015), and SimVerb-3500 Gerz et al. (2016). Each dataset comprises word pairs along with gold similarity scores, assigned by human annotators. We employed the Spearman rank correlation coefficient between human ratings and the cosine similarity of the word embeddings as the evaluation metric. The reported values represent the average scores across the six datasets.
Results and Discussion
The experimental results are presented in Table 5. We observe that ICA-based methods outperform PCA-based methods. Moreover, our proposed method, Spectral Clustering on the MST of ICA components, consistently achieves the best performance across all dimensions. This can be attributed to the fact that components included in the same cluster on the MST likely have high semantic relevance and play similar roles in representing the meaning of words. These results validate that considering higher-order correlations between axes better preserves semantic relationships in compressed word embeddings, demonstrating the practical utility of our method. It is important to note that this evaluation assesses the performance of clustering using a downstream task of dimensionality reduction, rather than dimensionality reduction itself.
Appendix G Supplementary Tables for ICA Components and MST Subtrees
Table 6 shows all components of the ICA-transformed word embeddings used in our experiments. Associated with the experiments in Sec. 4.1, Table 7 shows the top 60 pairs with the highest $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ values. Additionally, in Table 8 and Table 9, we report on all component pairs in the subtrees of MST $T_{150}$ shown in Fig. 5, extending the results from the selected pairs previously reported in Sec. 4.3.
Axis 00 Axis $1$ Axis $2$ Axis $3$ Axis $4$ Axis $5$ Axis $6$ Axis $7$ Axis $8$ Axis $9$ Axis $10$ Axis $11$ Axis $12$ Axis $13$ Axis $14$ dishes genus acid al india und el les chinese di dna gaius rabbi windows topological sauce species hydrogen ibn indian der spanish du china italian proteins caesar talmud os isomorphic fried extinct acids muhammad nehru die nacional des beijing della rna augustus rabbis unix banach dish birds oh abu hindu das jos paris pinyin luigi mrna lucius torah linux topology Axis $15$ Axis $16$ Axis $17$ Axis $18$ Axis $19$ Axis $20$ Axis $21$ Axis $22$ Axis $23$ Axis $24$ Axis $25$ Axis $26$ Axis $27$ Axis $28$ Axis $29$ drugs blood consonants actress party stars japanese instrument statesman church russian cars greek album film drug organs vowels footballer parties constellation japan instruments astronomer churches moscow ford greece albums films heroin liver vowel musician democrats star tokyo bass philosopher communion russia car athens band directed lsd kidney consonant actor democratic constellations emperor guitars johann orthodox soviet chassis athenian songs director Axis $30$ Axis $31$ Axis $32$ Axis $33$ Axis $34$ Axis $35$ Axis $36$ Axis $37$ Axis $38$ Axis $39$ Axis $40$ Axis $41$ Axis $42$ Axis $43$ Axis $44$ stations swedish university polish irish aircraft http rail rifle pitcher isbn islands jpg currency plants fm sweden college poland ireland flight www trains gun sox press island image currencies plant radio danish technology krak dublin boeing htm railway rifles baseman ed archipelago png euro flowers broadcast norwegian institute aw ulster airlines html train ammunition pitchers routledge atoll gif dollar flowering Axis $45$ Axis $46$ Axis $47$ Axis $48$ Axis $49$ Axis $50$ Axis $51$ Axis $52$ Axis $53$ Axis $54$ Axis $55$ Axis $56$ Axis $57$ Axis $58$ Axis $59$ quantum spacecraft ship day court site coach infectious element buildings appointed cpu s imperfect concerto particles nasa ships observances judge website quarterback infection metals building minister microprocessor and perfect fugue particle astronauts hms holidays courts forum defensive disease elements tower cabinet processor was future sonata physics astronaut cruisers holiday trial photos bengals infections uranium built appoints cpus in present bwv Axis $60$ Axis $61$ Axis $62$ Axis $63$ Axis $64$ Axis $65$ Axis $66$ Axis $67$ Axis $68$ Axis $69$ Axis $70$ Axis $71$ Axis $72$ Axis $73$ Axis $74$ river daughter company organization goddess word judah blue football gesserit accusative game sexual ip africa tributaries married corporation international gods meaning israelites white cup bene nouns games sex tcp african rivers marriage companies organizations deity etymology yahweh red fa leto genitive gameplay homosexual protocols africans navigable wife shareholders interpol deities term elisha yellow club duncan noun multiplayer heterosexual protocol namibia Axis $75$ Axis $76$ Axis $77$ Axis $78$ Axis $79$ Axis $80$ Axis $81$ Axis $82$ Axis $83$ Axis $84$ Axis $85$ Axis $86$ Axis $87$ Axis $88$ Axis $89$ births you painting fiction canada serbian philosophy beer wimbledon highest battle rocks solids force austria deaths know paintings novels quebec serbs philosophical beers open elevation battles volcanic faces army belgium alumni me art novel canadian serbia kant ale finalist ft defeat granite vertices military luxembourg novelists we painters stories ontario croatia philosophers brewing quarter lowest fought geologic cube regiment germany Axis $90$ Axis $91$ Axis $92$ Axis $93$ Axis $94$ Axis $95$ Axis $96$ Axis $97$ Axis $98$ Axis $99$ Axis $100$ Axis $101$ Axis $102$ Axis $103$ Axis $104$ comics administrative essex languages ethnic dutch o capitalism australia missouri empire voltage encryption wavelength archaeological marvel divided somerset spoken peoples van portuguese anarcho australian kentucky empires electrical cryptography light neolithic superhero districts cornwall language indigenous netherlands paulo capitalists sydney mississippi emperors circuits cipher wavelengths bc superman divisions exeter dialects minorities amsterdam rio economists melbourne alabama rulers current ciphers laser pottery Axis $105$ Axis $106$ Axis $107$ Axis $108$ Axis $109$ Axis $110$ Axis $111$ Axis $112$ Axis $113$ Axis $114$ Axis $115$ Axis $116$ Axis $117$ Axis $118$ Axis $119$ italy buddha dog award computation martial worn episode frac morphisms audio nazi si disorder combustion norway buddhism hound awards artificial judo wearing aired cos homomorphism lossy holocaust units mental diesel netherlands mahayana dogs prize turing aikido clothing show equation hydrogen compression nazis metre disorders turbine germany buddhist breed awarded computational karate wear tv euler wavelengths mpeg camps kilogram symptoms engine Axis $120$ Axis $121$ Axis $122$ Axis $123$ Axis $124$ Axis $125$ Axis $126$ Axis $127$ Axis $128$ Axis $129$ Axis $130$ Axis $131$ Axis $132$ Axis $133$ Axis $134$ winters families hip newspaper israeli algae pointer poetry ball schools exports iv telephone fourth prix summers family hop daily palestinian bacteria return verse scrimmage school imports iii phone third grand temperatures older dj weekly palestinians fungi string poem goal secondary textiles vii mobile fifth schumacher precipitation household rap newspapers israel mitochondria pointers poems foul education gwh vi cellular sixth race Axis $135$ Axis $136$ Axis $137$ Axis $138$ Axis $139$ Axis $140$ Axis $141$ Axis $142$ Axis $143$ Axis $144$ Axis $145$ Axis $146$ Axis $147$ Axis $148$ Axis $149$ manuscripts disk san horse frankish import treaty colspan card p donald studied medicine olympics saint translation floppy francisco horses franks duplicate signed align cards j duck born medical olympic st translations disks california riding saxon info agreement center dealer q scrooge he doctors athletes petersburg testament drives diego breed saxons no signing motto betting r mcduck career care sport patron Axis $150$ Axis $151$ Axis $152$ Axis $153$ Axis $154$ Axis $155$ Axis $156$ Axis $157$ Axis $158$ Axis $159$ Axis $160$ Axis $161$ Axis $162$ Axis $163$ Axis $164$ nuclear century inducted finns york list claim intelligence dance theatre evolution km ten corpus iran bomb th treasure soviets new topics claims agency dances stage evolutionary harbors hundred vitamin iranian bombs twentieth apocalypse col ny lists claimed cia dancing broadway darwin total months esp iraq fission nineteenth fairy ir bronx see evidence fbi dancers theater selection unpaved thousand mutant persian Axis $165$ Axis $166$ Axis $167$ Axis $168$ Axis $169$ Axis $170$ Axis $171$ Axis $172$ Axis $173$ Axis $174$ Axis $175$ Axis $176$ Axis $177$ Axis $178$ Axis $179$ desertification axis thank license female class monster hell basque park test article arthur question bah environment perpendicular hello copyleft male classes creature heaven spain parks tests main merlin debate ll flooding rotation please gpl age middle loch paradise aragon national testing disambiguation grail questions mason environmental direction good licenses infant working creatures eden eta wildlife cricket discusses knights whether meher Axis $180$ Axis $181$ Axis $182$ Axis $183$ Axis $184$ Axis $185$ Axis $186$ Axis $187$ Axis $188$ Axis $189$ Axis $190$ Axis $191$ Axis $192$ Axis $193$ Axis $194$ jedi coup ethiopia turkish g buried reading cort less taggart diamond melville guant rights freemasonry luke overthrow ethiopian turkey e burial further georges than pg diamonds fuller namo legislation masonic knight tat eritrea istanbul icao cemetery devised philip more estonian stones apr consistency act dawn wars junta kate cypriot fao grave steadily counts much finnish gem taste alter laws lodge Axis $195$ Axis $196$ Axis $197$ Axis $198$ Axis $199$ Axis $200$ Axis $201$ Axis $202$ Axis $203$ Axis $204$ Axis $205$ Axis $206$ Axis $207$ Axis $208$ Axis $209$ cretaceous indonesian population million derivative columbus actor cross centered maps names rolling chemistry short code geologic malaysia median estimated manchester taxi squadron crescent eagles map various adobe alchemy long calling epoch indonesia estimated billion resolutions benson actress icrc serial gis include chuck tubes lived codes extinction malaysian residing estimates pat concern isps dressing eventual gps these kasparov orientation conventional anthem Axis $210$ Axis $211$ Axis $212$ Axis $213$ Axis $214$ Axis $215$ Axis $216$ Axis $217$ Axis $218$ Axis $219$ Axis $220$ Axis $221$ Axis $222$ Axis $223$ Axis $224$ unesco calvinism k z tyler tex franklin esperanto tended foo cleese certificate oed anonymous caves itu laden defunct rich cornell marsh heath ido dave slight monty employee hungry gap alone interpol trotsky neutrality zwingli edit extermination falklands mi emissions needle python engineer angry collections elsewhere observer limited ln swiss theorem iliad boas ne assist folklore palin laplace threats incidence daniel Axis $225$ Axis $226$ Axis $227$ Axis $228$ Axis $229$ Axis $230$ Axis $231$ Axis $232$ Axis $233$ Axis $234$ Axis $235$ Axis $236$ Axis $237$ Axis $238$ Axis $239$ lord lenin principia pray bees spreading et aunt hong oswald pm voting mad martin pierre rings brands mainframe erasmus bee giving hoc blade kong bull tunnel plurality distinguishing luther lynx admiralty conan running raises ants popularity ad loan guidance op sec voter classify jr khmer courtesy russell stroke facing ant expertise est accidental mine distinct privileges voters karen zwingli bases Axis $240$ Axis $241$ Axis $242$ Axis $243$ Axis $244$ Axis $245$ Axis $246$ Axis $247$ Axis $248$ Axis $249$ Axis $250$ Axis $251$ Axis $252$ Axis $253$ Axis $254$ albert aether nor wood prevalence wayne guant sentences increase handling explosives ultimate joins jack advocacy einstein accelerated neither arrows worst ds namo lebesgue increased replication convenience reduction comparatively amos professor piercing populace anything aspartame cutting exhibits rn neil increasing oliver eccentric discover breed plug pirate combining bear exception merge bluetooth ears holes sigma increases ignore quasi striking fungi doug cuban Axis $255$ Axis $256$ Axis $257$ Axis $258$ Axis $259$ Axis $260$ Axis $261$ Axis $262$ Axis $263$ Axis $264$ Axis $265$ Axis $266$ Axis $267$ Axis $268$ Axis $269$ renowned pickford humanities fatal attending devastating cheeses accessed farmer metaphor landmarks glider textbook definitive criticized tourists coins statement provided xvi prints witchcraft arbor clown sanctioned cage trademark looks continually mead home turbulent conversation tcp fate assertion cheese ann desk strictly br purity springer eukaryotic stressed town earn witches descendant quad ink oath fahrenheit metaphysical sagan ruin flew seals gradient classified Axis $270$ Axis $271$ Axis $272$ Axis $273$ Axis $274$ Axis $275$ Axis $276$ Axis $277$ Axis $278$ Axis $279$ Axis $280$ Axis $281$ Axis $282$ Axis $283$ Axis $284$ postage doraemon importantly dodo crusader napoleon plots blake adjacent constructs feces cryonics cosmology economical multiply partly adaptation crusade girlfriend ceremony bonaparte hired ne proposed hierarchy buying ec reunited comprised guernsey willing inuit juan punk mailing herman twisted cap machines synod chess euthanasia ll initiative contraception binary bones carpenter mcgill euclidean reed characterize chances ban ois shadows freezing cosmological trap dependency Axis $285$ Axis $286$ Axis $287$ Axis $288$ Axis $289$ Axis $290$ Axis $291$ Axis $292$ Axis $293$ Axis $294$ Axis $295$ Axis $296$ Axis $297$ Axis $298$ Axis $299$ union wiki lynch mortal tributaries selling treated agave countryside beckham fitness fragile self cumberland gardner unrelated communities turned assassin sources copies desk expense accession analyses glass rousseau prince dylan correction lifelong usenet ordered fortified hormones aquatic sullivan audience burn pictures regular clever vested functions habsburg arm lightning meals dmt helicopters sell tables var completion trinidad marilyn planet ordained kay codes
Table 6: The top $4$ words with the largest component values along all axes of ICA-transformed word embedding used in our experiments. Axes are sorted in descending order of skewness.
$\mathrm{E}(S_{63}^{2}S_{210}^{2})=2.964$ $\mathrm{E}(S_{2}^{2}S_{114}^{2})=2.480$ $\mathrm{E}(S_{10}^{2}S_{125}^{2})=2.431$ $\mathrm{E}(S_{58}^{2}S_{70}^{2})=2.395$ $\mathrm{E}(S_{14}^{2}S_{113}^{2})=2.380$ $\mathrm{E}(S_{63}^{2}S_{184}^{2})=2.330$ Axis 63 Axis 210 Axis 2 Axis 114 Axis 10 Axis 125 Axis 58 Axis 70 Axis 14 Axis 113 Axis 63 Axis 184 organization unesco acid morphisms dna algae imperfect accusative topological frac organization g international itu hydrogen homomorphism proteins bacteria perfect nouns isomorphic cos international e organizations interpol acids hydrogen rna fungi future genitive banach equation organizations icao interpol observer oh wavelengths mrna mitochondria present noun topology euler interpol fao $\mathrm{E}(S_{2}^{2}S_{10}^{2})=2.323$ $\mathrm{E}(S_{22}^{2}S_{59}^{2})=2.247$ $\mathrm{E}(S_{26}^{2}S_{134}^{2})=2.233$ $\mathrm{E}(S_{14}^{2}S_{114}^{2})=2.228$ $\mathrm{E}(S_{2}^{2}S_{53}^{2})=2.165$ $\mathrm{E}(S_{16}^{2}S_{52}^{2})=2.158$ Axis 2 Axis 10 Axis 22 Axis 59 Axis 26 Axis 134 Axis 14 Axis 114 Axis 2 Axis 53 Axis 16 Axis 52 acid dna instrument concerto cars prix topological morphisms acid element blood infectious hydrogen proteins instruments fugue ford grand isomorphic homomorphism hydrogen metals organs infection acids rna bass sonata car schumacher banach hydrogen acids elements liver disease oh mrna guitars bwv chassis race topology wavelengths oh uranium kidney infections $\mathrm{E}(S_{10}^{2}S_{114}^{2})=2.150$ $\mathrm{E}(S_{16}^{2}S_{118}^{2})=2.124$ $\mathrm{E}(S_{1}^{2}S_{121}^{2})=2.107$ $\mathrm{E}(S_{44}^{2}S_{121}^{2})=2.096$ $\mathrm{E}(S_{28}^{2}S_{122}^{2})=2.086$ $\mathrm{E}(S_{19}^{2}S_{55}^{2})=2.075$ Axis 10 Axis 114 Axis 16 Axis 118 Axis 1 Axis 121 Axis 44 Axis 121 Axis 28 Axis 122 Axis 19 Axis 55 dna morphisms blood disorder genus families plants families album hip party appointed proteins homomorphism organs mental species family plant family albums hop parties minister rna hydrogen liver disorders extinct older flowers older band dj democrats cabinet mrna wavelengths kidney symptoms birds household flowering household songs rap democratic appoints $\mathrm{E}(S_{30}^{2}S_{112}^{2})=2.060$ $\mathrm{E}(S_{6}^{2}S_{96}^{2})=2.032$ $\mathrm{E}(S_{21}^{2}S_{110}^{2})=2.028$ $\mathrm{E}(S_{27}^{2}S_{64}^{2})=1.997$ $\mathrm{E}(S_{13}^{2}S_{168}^{2})=1.991$ $\mathrm{E}(S_{51}^{2}S_{128}^{2})=1.987$ Axis 30 Axis 112 Axis 6 Axis 96 Axis 21 Axis 110 Axis 27 Axis 64 Axis 13 Axis 168 Axis 51 Axis 128 stations episode el o japanese martial greek goddess windows license coach ball fm aired spanish portuguese japan judo greece gods os copyleft quarterback scrimmage radio show nacional paulo tokyo aikido athens deity unix gpl defensive goal broadcast tv jos rio emperor karate athenian deities linux licenses bengals foul $\mathrm{E}(S_{12}^{2}S_{66}^{2})=1.975$ $\mathrm{E}(S_{4}^{2}S_{106}^{2})=1.974$ $\mathrm{E}(S_{56}^{2}S_{136}^{2})=1.967$ $\mathrm{E}(S_{17}^{2}S_{114}^{2})=1.966$ $\mathrm{E}(S_{15}^{2}S_{118}^{2})=1.959$ $\mathrm{E}(S_{10}^{2}S_{16}^{2})=1.947$ Axis 12 Axis 66 Axis 4 Axis 106 Axis 56 Axis 136 Axis 17 Axis 114 Axis 15 Axis 118 Axis 10 Axis 16 rabbi judah india buddha cpu disk consonants morphisms drugs disorder dna blood talmud israelites indian buddhism microprocessor floppy vowels homomorphism drug mental proteins organs rabbis yahweh nehru mahayana processor disks vowel hydrogen heroin disorders rna liver torah elisha hindu buddhist cpus drives consonant wavelengths lsd symptoms mrna kidney $\mathrm{E}(S_{0}^{2}S_{82}^{2})=1.927$ $\mathrm{E}(S_{53}^{2}S_{150}^{2})=1.915$ $\mathrm{E}(S_{80}^{2}S_{89}^{2})=1.909$ $\mathrm{E}(S_{5}^{2}S_{59}^{2})=1.897$ $\mathrm{E}(S_{39}^{2}S_{51}^{2})=1.885$ $\mathrm{E}(S_{10}^{2}S_{52}^{2})=1.875$ Axis 0 Axis 82 Axis 53 Axis 150 Axis 80 Axis 89 Axis 5 Axis 59 Axis 39 Axis 51 Axis 10 Axis 52 dishes beer element nuclear serbian austria und concerto pitcher coach dna infectious sauce beers metals bomb serbs belgium der fugue sox quarterback proteins infection fried ale elements bombs serbia luxembourg die sonata baseman defensive rna disease dish brewing uranium fission croatia germany das bwv pitchers bengals mrna infections $\mathrm{E}(S_{53}^{2}S_{86}^{2})=1.872$ $\mathrm{E}(S_{39}^{2}S_{128}^{2})=1.871$ $\mathrm{E}(S_{6}^{2}S_{173}^{2})=1.869$ $\mathrm{E}(S_{8}^{2}S_{106}^{2})=1.867$ $\mathrm{E}(S_{56}^{2}S_{126}^{2})=1.861$ $\mathrm{E}(S_{85}^{2}S_{153}^{2})=1.857$ Axis 53 Axis 86 Axis 39 Axis 128 Axis 6 Axis 173 Axis 8 Axis 106 Axis 56 Axis 126 Axis 85 Axis 153 element rocks pitcher ball el basque chinese buddha cpu pointer battle finns metals volcanic sox scrimmage spanish spain china buddhism microprocessor return battles soviets elements granite baseman goal nacional aragon beijing mahayana processor string defeat col uranium geologic pitchers foul jos eta pinyin buddhist cpus pointers fought ir $\mathrm{E}(S_{74}^{2}S_{89}^{2})=1.854$ $\mathrm{E}(S_{15}^{2}S_{16}^{2})=1.834$ $\mathrm{E}(S_{83}^{2}S_{105}^{2})=1.832$ $\mathrm{E}(S_{17}^{2}S_{233}^{2})=1.828$ $\mathrm{E}(S_{71}^{2}S_{143}^{2})=1.828$ $\mathrm{E}(S_{53}^{2}S_{114}^{2})=1.824$ Axis 74 Axis 89 Axis 15 Axis 16 Axis 83 Axis 105 Axis 17 Axis 233 Axis 71 Axis 143 Axis 53 Axis 114 africa austria drugs blood wimbledon italy consonants hong game card element morphisms african belgium drug organs open norway vowels kong games cards metals homomorphism africans luxembourg heroin liver finalist netherlands vowel guidance gameplay dealer elements hydrogen namibia germany lsd kidney quarter germany consonant mine multiplayer betting uranium wavelengths $\mathrm{E}(S_{30}^{2}S_{132}^{2})=1.820$ $\mathrm{E}(S_{10}^{2}S_{160}^{2})=1.811$ $\mathrm{E}(S_{86}^{2}S_{195}^{2})=1.809$ $\mathrm{E}(S_{101}^{2}S_{103}^{2})=1.800$ $\mathrm{E}(S_{35}^{2}S_{88}^{2})=1.798$ $\mathrm{E}(S_{4}^{2}S_{89}^{2})=1.794$ Axis 30 Axis 132 Axis 10 Axis 160 Axis 86 Axis 195 Axis 101 Axis 103 Axis 35 Axis 88 Axis 4 Axis 89 stations telephone dna evolution rocks cretaceous voltage wavelength aircraft force india austria fm phone proteins evolutionary volcanic geologic electrical light flight army indian belgium radio mobile rna darwin granite epoch circuits wavelengths boeing military nehru luxembourg broadcast cellular mrna selection geologic extinction current laser airlines regiment hindu germany $\mathrm{E}(S_{107}^{2}S_{138}^{2})=1.793$ $\mathrm{E}(S_{15}^{2}S_{44}^{2})=1.793$ $\mathrm{E}(S_{32}^{2}S_{129}^{2})=1.793$ $\mathrm{E}(S_{20}^{2}S_{46}^{2})=1.788$ $\mathrm{E}(S_{3}^{2}S_{124}^{2})=1.786$ $\mathrm{E}(S_{26}^{2}S_{119}^{2})=1.785$ Axis 107 Axis 138 Axis 15 Axis 44 Axis 32 Axis 129 Axis 20 Axis 46 Axis 3 Axis 124 Axis 26 Axis 119 dog horse drugs plants university schools stars spacecraft al israeli cars combustion hound horses drug plant college school constellation nasa ibn palestinian ford diesel dogs riding heroin flowers technology secondary star astronauts muhammad palestinians car turbine breed breed lsd flowering institute education constellations astronaut abu israel chassis engine
Table 7: Complementary experimental results to Table 1. The top 60 pairs with the highest $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ values are presented. For each component, the top 4 words with the largest component values are listed.
$\mathrm{E}(S_{2}^{2}S_{10}^{2})=2.323$ $\mathrm{E}(S_{2}^{2}S_{119}^{2})=1.755$ $\mathrm{E}(S_{132}^{2}S_{30}^{2})=1.820$ $\mathrm{E}(S_{132}^{2}S_{73}^{2})=1.693$ $\mathrm{E}(S_{132}^{2}S_{62}^{2})=1.632$ $\mathrm{E}(S_{136}^{2}S_{56}^{2})=1.967$ Axis $2$ Axis $10$ Axis $2$ Axis $119$ Axis $132$ Axis $30$ Axis $132$ Axis $73$ Axis $132$ Axis $62$ Axis $136$ Axis $56$ acid dna acid combustion telephone stations telephone ip telephone company disk cpu hydrogen proteins hydrogen diesel phone fm phone tcp phone corporation floppy microprocessor acids rna acids turbine mobile radio mobile protocols mobile companies disks processor oh mrna oh engine cellular broadcast cellular protocol cellular shareholders drives cpus $w_{k}$ $\mathbf{S}_{k,2}^{2}\mathbf{S}_{k,{10}}^{2}$ $w_{k}$ $\mathbf{S}_{k,2}^{2}\mathbf{S}_{k,{119}}^{2}$ $w_{k}$ $\mathbf{S}_{k,132}^{2}\mathbf{S}_{k,{30}}^{2}$ $w_{k}$ $\mathbf{S}_{k,132}^{2}\mathbf{S}_{k,{73}}^{2}$ $w_{k}$ $\mathbf{S}_{k,132}^{2}\mathbf{S}_{k,{62}}^{2}$ $w_{k}$ $\mathbf{S}_{k,136}^{2}\mathbf{S}_{k,{56}}^{2}$ ribose 3755.7 pyrolysis 2794.4 digitalized 4726.8 multipoint 2062.7 esat 3145.5 sata 3427.5 deoxyribose 2963.9 syngas 2056.9 arabsat 4657.2 pstn 1996.9 telecoms 2547.9 udma 2519.7 phosphodiester 2850.2 gasification 1783.5 radiotelephone 4453.9 wimax 1873.8 nynex 2155.9 backplanes 2008.9 biosynthesis 2510.1 butane 1761.2 landlines 3522.1 xdsl 1491.2 gnc 1810.0 nexgen 1947.7 methyltransferase 2482.9 dehydrogenation 1623.0 intersputnik 3053.5 svcs 1361.2 haitel 1657.3 megabytes 1890.1 pyrimidine 2399.6 tert 1230.2 telex 2722.4 isdn 1235.8 openreach 1529.2 eisa 1859.5 $\mathrm{E}(S_{10}^{2}S_{16}^{2})=1.947$ $\mathrm{E}(S_{10}^{2}S_{160}^{2})=1.811$ $\mathrm{E}(S_{140}^{2}S_{56}^{2})=1.615$ $\mathrm{E}(S_{13}^{2}S_{56}^{2})=1.732$ $\mathrm{E}(S_{13}^{2}S_{168}^{2})=1.991$ $\mathrm{E}(S_{13}^{2}S_{73}^{2})=1.740$ Axis $10$ Axis $16$ Axis $10$ Axis $160$ Axis $140$ Axis $56$ Axis $13$ Axis $56$ Axis $13$ Axis $168$ Axis $13$ Axis $73$ dna blood dna evolution import cpu windows cpu windows license windows ip proteins organs proteins evolutionary duplicate microprocessor os microprocessor os copyleft os tcp rna liver rna darwin info processor unix processor unix gpl unix protocols mrna kidney mrna selection no cpus linux cpus linux licenses linux protocol $w_{k}$ $\mathbf{S}_{k,10}^{2}\mathbf{S}_{k,{16}}^{2}$ $w_{k}$ $\mathbf{S}_{k,10}^{2}\mathbf{S}_{k,{160}}^{2}$ $w_{k}$ $\mathbf{S}_{k,140}^{2}\mathbf{S}_{k,{56}}^{2}$ $w_{k}$ $\mathbf{S}_{k,13}^{2}\mathbf{S}_{k,{56}}^{2}$ $w_{k}$ $\mathbf{S}_{k,13}^{2}\mathbf{S}_{k,{168}}^{2}$ $w_{k}$ $\mathbf{S}_{k,13}^{2}\mathbf{S}_{k,{73}}^{2}$ adenylate 2079.8 utr 2381.5 superpipelined 5652.9 xcode 2609.7 qpl 5678.2 netware 1799.2 effectors 1842.5 reticulum 1942.0 strongarm 3220.2 powerpc 2046.7 lgpl 4519.9 netbios 1543.2 antisense 1639.9 genomic 1668.6 specrate 2470.5 itanium 1500.1 trolltech 3588.4 imap 1414.0 cyclase 1638.9 homozygous 1599.1 specbaserate 1524.5 glibc 1355.1 gpl 3325.2 glut 1239.0 myosin 1201.8 cleaved 1181.0 insubstantial 1387.8 irix 1177.1 gnu 2826.1 wfw 1179.9 axons 1144.2 tubulin 1152.4 eisa 1027.8 efi 1161.1 bsd 2822.7 dhcpv 1115.5 $\mathrm{E}(S_{15}^{2}S_{118}^{2})=1.959$ $\mathrm{E}(S_{15}^{2}S_{44}^{2})=1.793$ $\mathrm{E}(S_{16}^{2}S_{248}^{2})=1.499$ $\mathrm{E}(S_{16}^{2}S_{147}^{2})=1.717$ $\mathrm{E}(S_{16}^{2}S_{72}^{2})=1.628$ $\mathrm{E}(S_{16}^{2}S_{52}^{2})=2.158$ Axis $15$ Axis $118$ Axis $15$ Axis $44$ Axis $16$ Axis $248$ Axis $16$ Axis $147$ Axis $16$ Axis $72$ Axis $16$ Axis $52$ drugs disorder drugs plants blood increase blood medicine blood sexual blood infectious drug mental drug plant organs increased organs medical organs sex organs infection heroin disorders heroin flowers liver increasing liver doctors liver homosexual liver disease lsd symptoms lsd flowering kidney increases kidney care kidney heterosexual kidney infections $w_{k}$ $\mathbf{S}_{k,15}^{2}\mathbf{S}_{k,{118}}^{2}$ $w_{k}$ $\mathbf{S}_{k,15}^{2}\mathbf{S}_{k,{44}}^{2}$ $w_{k}$ $\mathbf{S}_{k,16}^{2}\mathbf{S}_{k,{248}}^{2}$ $w_{k}$ $\mathbf{S}_{k,16}^{2}\mathbf{S}_{k,{147}}^{2}$ $w_{k}$ $\mathbf{S}_{k,16}^{2}\mathbf{S}_{k,{72}}^{2}$ $w_{k}$ $\mathbf{S}_{k,16}^{2}\mathbf{S}_{k,{52}}^{2}$ adhd 3505.8 peyote 3197.5 esophagus 2274.6 aortic 1726.9 erectile 2001.5 abscess 1932.6 anticonvulsants 3047.9 purpurea 2926.6 lobes 1462.4 brainstem 1537.1 dildo 1788.8 multifocal 1440.2 sertraline 2374.6 meo 2878.6 cava 1360.8 endoscopy 1533.7 clitoral 1481.5 hemorrhagic 1239.3 lorazepam 1604.8 deliriants 2397.0 ligamentum 1024.8 laparoscopic 1509.4 deferens 1432.6 esophagitis 1187.4 anticonvulsant 1487.1 diplopterys 2395.1 transversal 1011.2 angioplasty 1447.4 rectal 1352.0 efferent 1160.5 somnolence 1426.1 cabrerana 2237.7 vena 1005.6 cardiology 1433.3 urogenital 1344.7 mitral 1143.5 $\mathrm{E}(S_{16}^{2}S_{118}^{2})=2.124$ $\mathrm{E}(S_{30}^{2}S_{112}^{2})=2.060$ $\mathrm{E}(S_{30}^{2}S_{123}^{2})=1.570$ $\mathrm{E}(S_{36}^{2}S_{50}^{2})=1.688$ $\mathrm{E}(S_{36}^{2}S_{73}^{2})=1.419$ $\mathrm{E}(S_{168}^{2}S_{193}^{2})=1.495$ Axis $16$ Axis $118$ Axis $30$ Axis $112$ Axis $30$ Axis $123$ Axis $36$ Axis $50$ Axis $36$ Axis $73$ Axis $168$ Axis $193$ blood disorder stations episode stations newspaper http site http ip license rights organs mental fm aired fm daily www website www tcp copyleft legislation liver disorders radio show radio weekly htm forum htm protocols gpl act kidney symptoms broadcast tv broadcast newspapers html photos html protocol licenses laws $w_{k}$ $\mathbf{S}_{k,16}^{2}\mathbf{S}_{k,{118}}^{2}$ $w_{k}$ $\mathbf{S}_{k,30}^{2}\mathbf{S}_{k,{112}}^{2}$ $w_{k}$ $\mathbf{S}_{k,30}^{2}\mathbf{S}_{k,{123}}^{2}$ $w_{k}$ $\mathbf{S}_{k,36}^{2}\mathbf{S}_{k,{50}}^{2}$ $w_{k}$ $\mathbf{S}_{k,36}^{2}\mathbf{S}_{k,{73}}^{2}$ $w_{k}$ $\mathbf{S}_{k,168}^{2}\mathbf{S}_{k,{193}}^{2}$ atrophy 2110.2 rebroadcast 1729.4 canwest 1941.7 shtml 1537.9 mtu 1425.4 copyleft 1648.0 hemiparesis 1877.5 fsn 1635.4 ctv 1343.4 geocities 1230.3 stateful 1352.4 magnatune 1131.9 axonal 1465.9 etv 1600.9 wqxr 1276.0 jeancocteau 871.2 proxying 1324.2 rightsholder 839.7 dysfunction 1380.2 upn 1534.6 superstation 1144.2 lfc 758.0 mpls 1264.5 redistribute 809.7 neuropathy 1300.1 wxyz 1441.3 wanbao 1116.1 uchicago 644.7 vpns 798.6 copyrights 676.5 myopathy 1288.3 whdh 1392.0 aor 998.0 artchive 593.3 kleinrock 796.1 circumvention 653.3 $\mathrm{E}(S_{169}^{2}S_{72}^{2})=1.701$ $\mathrm{E}(S_{44}^{2}S_{121}^{2})=2.096$ $\mathrm{E}(S_{45}^{2}S_{103}^{2})=1.765$ $\mathrm{E}(S_{49}^{2}S_{193}^{2})=1.616$ $\mathrm{E}(S_{56}^{2}S_{101}^{2})=1.693$ $\mathrm{E}(S_{56}^{2}S_{71}^{2})=1.592$ Axis $169$ Axis $72$ Axis $44$ Axis $121$ Axis $45$ Axis $103$ Axis $49$ Axis $193$ Axis $56$ Axis $101$ Axis $56$ Axis $71$ female sexual plants families quantum wavelength court rights cpu voltage cpu game male sex plant family particles light judge legislation microprocessor electrical microprocessor games age homosexual flowers older particle wavelengths courts act processor circuits processor gameplay infant heterosexual flowering household physics laser trial laws cpus current cpus multiplayer $w_{k}$ $\mathbf{S}_{k,169}^{2}\mathbf{S}_{k,{72}}^{2}$ $w_{k}$ $\mathbf{S}_{k,44}^{2}\mathbf{S}_{k,{121}}^{2}$ $w_{k}$ $\mathbf{S}_{k,45}^{2}\mathbf{S}_{k,{103}}^{2}$ $w_{k}$ $\mathbf{S}_{k,49}^{2}\mathbf{S}_{k,{193}}^{2}$ $w_{k}$ $\mathbf{S}_{k,56}^{2}\mathbf{S}_{k,{101}}^{2}$ $w_{k}$ $\mathbf{S}_{k,56}^{2}\mathbf{S}_{k,{71}}^{2}$ male 1240.5 rosid 4158.0 mesons 3078.7 habeas 954.7 lsi 2589.8 backlit 1398.9 vulval 1191.4 caryophyllales 4064.1 gluons 2079.4 declaratory 823.5 microelectronic 2081.5 quadra 1388.5 faggot 981.2 dicotyledons 3652.5 photon 1648.3 conservatorship 783.1 sram 1901.2 gba 1349.0 spermatozoon 961.5 betulaceae 3637.6 photons 1567.6 waives 753.2 mosfet 1757.0 vectrex 1128.1 frot 940.6 cronquist 3636.1 synchrotron 1435.1 laches 734.4 voltages 1582.9 epyx 1113.6 tribadism 843.5 poaceae 3295.0 isospin 1414.1 talionis 721.3 microcontrollers 1554.5 vcs 1097.0 $\mathrm{E}(S_{188}^{2}S_{119}^{2})=1.442$ $\mathrm{E}(S_{188}^{2}S_{205}^{2})=1.306$ $\mathrm{E}(S_{67}^{2}S_{111}^{2})=1.723$ $\mathrm{E}(S_{67}^{2}S_{103}^{2})=1.705$ $\mathrm{E}(S_{101}^{2}S_{103}^{2})=1.800$ $\mathrm{E}(S_{101}^{2}S_{119}^{2})=1.641$ Axis $188$ Axis $119$ Axis $188$ Axis $205$ Axis $67$ Axis $111$ Axis $67$ Axis $103$ Axis $101$ Axis $103$ Axis $101$ Axis $119$ less combustion less names blue worn blue wavelength voltage wavelength voltage combustion than diesel than various white wearing white light electrical light electrical diesel more turbine more include red clothing red wavelengths circuits wavelengths circuits turbine much engine much these yellow wear yellow laser current laser current engine $w_{k}$ $\mathbf{S}_{k,188}^{2}\mathbf{S}_{k,{119}}^{2}$ $w_{k}$ $\mathbf{S}_{k,188}^{2}\mathbf{S}_{k,{205}}^{2}$ $w_{k}$ $\mathbf{S}_{k,67}^{2}\mathbf{S}_{k,{111}}^{2}$ $w_{k}$ $\mathbf{S}_{k,67}^{2}\mathbf{S}_{k,{103}}^{2}$ $w_{k}$ $\mathbf{S}_{k,101}^{2}\mathbf{S}_{k,{103}}^{2}$ $w_{k}$ $\mathbf{S}_{k,101}^{2}\mathbf{S}_{k,{119}}^{2}$ thermojet 816.0 gec 505.9 sash 2800.1 thz 2245.4 photodiode 2784.6 alternator 1508.5 dirtier 805.6 nervosa 481.3 hakama 2144.4 monochromatic 1777.2 diodes 2320.6 cogeneration 1485.3 cng 739.9 dddddd 430.1 turbans 1756.4 protanomaly 1545.8 diode 1959.3 clamp 1370.2 oxidiser 655.3 ler 373.6 waists 1745.6 fluoresce 1376.9 ssi 1633.5 solenoid 1183.5 tuyere 634.7 gomoku 356.8 tunics 1612.8 scintillator 1366.4 bandgap 1523.4 thermionic 1144.8 propfan 630.0 kel 317.6 pleated 1561.4 diffuser 1233.8 photodiodes 1478.2 transformer 1055.4
Table 8: Complementary experimental results to Table 3. For all component pairs $(S_{i},S_{j})$ in the first subtree of the MST in Fig. 5, the top 6 words and their corresponding $\mathbf{S}_{t,i}^{2}\mathbf{S}_{t,j}^{2}$ values that contribute the most to the $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ value are presented.
$\mathrm{E}(S_{2}^{2}S_{10}^{2})=2.323$ $\mathrm{E}(S_{2}^{2}S_{119}^{2})=1.755$ $\mathrm{E}(S_{132}^{2}S_{30}^{2})=1.820$ $\mathrm{E}(S_{132}^{2}S_{73}^{2})=1.693$ $\mathrm{E}(S_{132}^{2}S_{62}^{2})=1.632$ $\mathrm{E}(S_{136}^{2}S_{56}^{2})=1.967$ Axis $2$ Axis $10$ Axis $2$ Axis $119$ Axis $132$ Axis $30$ Axis $132$ Axis $73$ Axis $132$ Axis $62$ Axis $136$ Axis $56$ acid dna acid combustion telephone stations telephone ip telephone company disk cpu hydrogen proteins hydrogen diesel phone fm phone tcp phone corporation floppy microprocessor acids rna acids turbine mobile radio mobile protocols mobile companies disks processor oh mrna oh engine cellular broadcast cellular protocol cellular shareholders drives cpus $w_{k}$ $\mathbf{S}_{k,2}^{2}\mathbf{S}_{k,{10}}^{2}$ $w_{k}$ $\mathbf{S}_{k,2}^{2}\mathbf{S}_{k,{119}}^{2}$ $w_{k}$ $\mathbf{S}_{k,132}^{2}\mathbf{S}_{k,{30}}^{2}$ $w_{k}$ $\mathbf{S}_{k,132}^{2}\mathbf{S}_{k,{73}}^{2}$ $w_{k}$ $\mathbf{S}_{k,132}^{2}\mathbf{S}_{k,{62}}^{2}$ $w_{k}$ $\mathbf{S}_{k,136}^{2}\mathbf{S}_{k,{56}}^{2}$ ribose 3755.7 pyrolysis 2794.4 digitalized 4726.8 multipoint 2062.7 esat 3145.5 sata 3427.5 deoxyribose 2963.9 syngas 2056.9 arabsat 4657.2 pstn 1996.9 telecoms 2547.9 udma 2519.7 phosphodiester 2850.2 gasification 1783.5 radiotelephone 4453.9 wimax 1873.8 nynex 2155.9 backplanes 2008.9 biosynthesis 2510.1 butane 1761.2 landlines 3522.1 xdsl 1491.2 gnc 1810.0 nexgen 1947.7 methyltransferase 2482.9 dehydrogenation 1623.0 intersputnik 3053.5 svcs 1361.2 haitel 1657.3 megabytes 1890.1 pyrimidine 2399.6 tert 1230.2 telex 2722.4 isdn 1235.8 openreach 1529.2 eisa 1859.5 $\mathrm{E}(S_{10}^{2}S_{16}^{2})=1.947$ $\mathrm{E}(S_{10}^{2}S_{160}^{2})=1.811$ $\mathrm{E}(S_{140}^{2}S_{56}^{2})=1.615$ $\mathrm{E}(S_{13}^{2}S_{56}^{2})=1.732$ $\mathrm{E}(S_{13}^{2}S_{168}^{2})=1.991$ $\mathrm{E}(S_{13}^{2}S_{73}^{2})=1.740$ Axis $10$ Axis $16$ Axis $10$ Axis $160$ Axis $140$ Axis $56$ Axis $13$ Axis $56$ Axis $13$ Axis $168$ Axis $13$ Axis $73$ dna blood dna evolution import cpu windows cpu windows license windows ip proteins organs proteins evolutionary duplicate microprocessor os microprocessor os copyleft os tcp rna liver rna darwin info processor unix processor unix gpl unix protocols mrna kidney mrna selection no cpus linux cpus linux licenses linux protocol $w_{k}$ $\mathbf{S}_{k,10}^{2}\mathbf{S}_{k,{16}}^{2}$ $w_{k}$ $\mathbf{S}_{k,10}^{2}\mathbf{S}_{k,{160}}^{2}$ $w_{k}$ $\mathbf{S}_{k,140}^{2}\mathbf{S}_{k,{56}}^{2}$ $w_{k}$ $\mathbf{S}_{k,13}^{2}\mathbf{S}_{k,{56}}^{2}$ $w_{k}$ $\mathbf{S}_{k,13}^{2}\mathbf{S}_{k,{168}}^{2}$ $w_{k}$ $\mathbf{S}_{k,13}^{2}\mathbf{S}_{k,{73}}^{2}$ adenylate 2079.8 utr 2381.5 superpipelined 5652.9 xcode 2609.7 qpl 5678.2 netware 1799.2 effectors 1842.5 reticulum 1942.0 strongarm 3220.2 powerpc 2046.7 lgpl 4519.9 netbios 1543.2 antisense 1639.9 genomic 1668.6 specrate 2470.5 itanium 1500.1 trolltech 3588.4 imap 1414.0 cyclase 1638.9 homozygous 1599.1 specbaserate 1524.5 glibc 1355.1 gpl 3325.2 glut 1239.0 myosin 1201.8 cleaved 1181.0 insubstantial 1387.8 irix 1177.1 gnu 2826.1 wfw 1179.9 axons 1144.2 tubulin 1152.4 eisa 1027.8 efi 1161.1 bsd 2822.7 dhcpv 1115.5 $\mathrm{E}(S_{15}^{2}S_{118}^{2})=1.959$ $\mathrm{E}(S_{15}^{2}S_{44}^{2})=1.793$ $\mathrm{E}(S_{16}^{2}S_{248}^{2})=1.499$ $\mathrm{E}(S_{16}^{2}S_{147}^{2})=1.717$ $\mathrm{E}(S_{16}^{2}S_{72}^{2})=1.628$ $\mathrm{E}(S_{16}^{2}S_{52}^{2})=2.158$ Axis $15$ Axis $118$ Axis $15$ Axis $44$ Axis $16$ Axis $248$ Axis $16$ Axis $147$ Axis $16$ Axis $72$ Axis $16$ Axis $52$ drugs disorder drugs plants blood increase blood medicine blood sexual blood infectious drug mental drug plant organs increased organs medical organs sex organs infection heroin disorders heroin flowers liver increasing liver doctors liver homosexual liver disease lsd symptoms lsd flowering kidney increases kidney care kidney heterosexual kidney infections $w_{k}$ $\mathbf{S}_{k,15}^{2}\mathbf{S}_{k,{118}}^{2}$ $w_{k}$ $\mathbf{S}_{k,15}^{2}\mathbf{S}_{k,{44}}^{2}$ $w_{k}$ $\mathbf{S}_{k,16}^{2}\mathbf{S}_{k,{248}}^{2}$ $w_{k}$ $\mathbf{S}_{k,16}^{2}\mathbf{S}_{k,{147}}^{2}$ $w_{k}$ $\mathbf{S}_{k,16}^{2}\mathbf{S}_{k,{72}}^{2}$ $w_{k}$ $\mathbf{S}_{k,16}^{2}\mathbf{S}_{k,{52}}^{2}$ adhd 3505.8 peyote 3197.5 esophagus 2274.6 aortic 1726.9 erectile 2001.5 abscess 1932.6 anticonvulsants 3047.9 purpurea 2926.6 lobes 1462.4 brainstem 1537.1 dildo 1788.8 multifocal 1440.2 sertraline 2374.6 meo 2878.6 cava 1360.8 endoscopy 1533.7 clitoral 1481.5 hemorrhagic 1239.3 lorazepam 1604.8 deliriants 2397.0 ligamentum 1024.8 laparoscopic 1509.4 deferens 1432.6 esophagitis 1187.4 anticonvulsant 1487.1 diplopterys 2395.1 transversal 1011.2 angioplasty 1447.4 rectal 1352.0 efferent 1160.5 somnolence 1426.1 cabrerana 2237.7 vena 1005.6 cardiology 1433.3 urogenital 1344.7 mitral 1143.5 $\mathrm{E}(S_{16}^{2}S_{118}^{2})=2.124$ $\mathrm{E}(S_{30}^{2}S_{112}^{2})=2.060$ $\mathrm{E}(S_{30}^{2}S_{123}^{2})=1.570$ $\mathrm{E}(S_{36}^{2}S_{50}^{2})=1.688$ $\mathrm{E}(S_{36}^{2}S_{73}^{2})=1.419$ $\mathrm{E}(S_{168}^{2}S_{193}^{2})=1.495$ Axis $16$ Axis $118$ Axis $30$ Axis $112$ Axis $30$ Axis $123$ Axis $36$ Axis $50$ Axis $36$ Axis $73$ Axis $168$ Axis $193$ blood disorder stations episode stations newspaper http site http ip license rights organs mental fm aired fm daily www website www tcp copyleft legislation liver disorders radio show radio weekly htm forum htm protocols gpl act kidney symptoms broadcast tv broadcast newspapers html photos html protocol licenses laws $w_{k}$ $\mathbf{S}_{k,16}^{2}\mathbf{S}_{k,{118}}^{2}$ $w_{k}$ $\mathbf{S}_{k,30}^{2}\mathbf{S}_{k,{112}}^{2}$ $w_{k}$ $\mathbf{S}_{k,30}^{2}\mathbf{S}_{k,{123}}^{2}$ $w_{k}$ $\mathbf{S}_{k,36}^{2}\mathbf{S}_{k,{50}}^{2}$ $w_{k}$ $\mathbf{S}_{k,36}^{2}\mathbf{S}_{k,{73}}^{2}$ $w_{k}$ $\mathbf{S}_{k,168}^{2}\mathbf{S}_{k,{193}}^{2}$ atrophy 2110.2 rebroadcast 1729.4 canwest 1941.7 shtml 1537.9 mtu 1425.4 copyleft 1648.0 hemiparesis 1877.5 fsn 1635.4 ctv 1343.4 geocities 1230.3 stateful 1352.4 magnatune 1131.9 axonal 1465.9 etv 1600.9 wqxr 1276.0 jeancocteau 871.2 proxying 1324.2 rightsholder 839.7 dysfunction 1380.2 upn 1534.6 superstation 1144.2 lfc 758.0 mpls 1264.5 redistribute 809.7 neuropathy 1300.1 wxyz 1441.3 wanbao 1116.1 uchicago 644.7 vpns 798.6 copyrights 676.5 myopathy 1288.3 whdh 1392.0 aor 998.0 artchive 593.3 kleinrock 796.1 circumvention 653.3 $\mathrm{E}(S_{169}^{2}S_{72}^{2})=1.701$ $\mathrm{E}(S_{44}^{2}S_{121}^{2})=2.096$ $\mathrm{E}(S_{45}^{2}S_{103}^{2})=1.765$ $\mathrm{E}(S_{49}^{2}S_{193}^{2})=1.616$ $\mathrm{E}(S_{56}^{2}S_{101}^{2})=1.693$ Axis $169$ Axis $72$ Axis $44$ Axis $121$ Axis $45$ Axis $103$ Axis $49$ Axis $193$ Axis $56$ Axis $101$ female sexual plants families quantum wavelength court rights cpu voltage male sex plant family particles light judge legislation microprocessor electrical age homosexual flowers older particle wavelengths courts act processor circuits infant heterosexual flowering household physics laser trial laws cpus current $w_{k}$ $\mathbf{S}_{k,169}^{2}\mathbf{S}_{k,{72}}^{2}$ $w_{k}$ $\mathbf{S}_{k,44}^{2}\mathbf{S}_{k,{121}}^{2}$ $w_{k}$ $\mathbf{S}_{k,45}^{2}\mathbf{S}_{k,{103}}^{2}$ $w_{k}$ $\mathbf{S}_{k,49}^{2}\mathbf{S}_{k,{193}}^{2}$ $w_{k}$ $\mathbf{S}_{k,56}^{2}\mathbf{S}_{k,{101}}^{2}$ male 1240.5 rosid 4158.0 mesons 3078.7 habeas 954.7 lsi 2589.8 vulval 1191.4 caryophyllales 4064.1 gluons 2079.4 declaratory 823.5 microelectronic 2081.5 faggot 981.2 dicotyledons 3652.5 photon 1648.3 conservatorship 783.1 sram 1901.2 spermatozoon 961.5 betulaceae 3637.6 photons 1567.6 waives 753.2 mosfet 1757.0 frot 940.6 cronquist 3636.1 synchrotron 1435.1 laches 734.4 voltages 1582.9 tribadism 843.5 poaceae 3295.0 isospin 1414.1 talionis 721.3 microcontrollers 1554.5
Table 9: Complementary experimental results to Table 3. For all component pairs $(S_{i},S_{j})$ in the second subtree of the MST in Fig. 5, the top 6 words and their corresponding $\mathbf{S}_{t,i}^{2}\mathbf{S}_{t,j}^{2}$ values that contribute the most to the $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ value are presented.